# PartSDF: Part-Based Implicit Neural Representation for Composite 3D Shape Parametrization and Optimization

### v1

@@ Abstract
Accurate 3D shape representation is essential in engineering applications such as design, optimization, and simulation. In practice, engineering workflows require structured, part-aware representations, as objects are inherently designed as assemblies of distinct components. However, most existing methods either model shapes holistically or decompose them without predefined part structures, limiting their applicability in real-world design tasks. We propose PartSDF, a supervised implicit representation framework that explicitly models composite shapes with independent, controllable parts while maintaining shape consistency. Despite its simple single-decoder architecture, PartSDF outperforms both supervised and unsupervised baselines in reconstruction and generation tasks. We further demonstrate its effectiveness as a structured shape prior for engineering applications, enabling precise control over individual components while preserving overall coherence. Code available at https://github.com/cvlab-epfl/PartSDF.
3d shape representation, INR, composite shape, shape optimization
@@ 1 Introduction
Figure 1: Part-based implicit representation. PartSDF is a modular approach to representing composite 3D shapes for a variety of purposes thanks to the simplicity of our part decoder: (a) shape generation, optionally conditioned on part layouts, and part-based (b) manipulation and (c) optimization. (c) The car body (in red) is optimized to reduce aerodynamic drag, without editing the wheels.
We live in a world full of manufactured objects of ever-increasing complexity that require clever engineering to be functional. These objects are inherently composite. Thus Computer Aided Design (CAD) tools used to design and optimize them must be able to handle them as assemblies of distinct parts from the outset. Each part must adhere to separate fabrication constraints, tolerances, and functional requirements, making part-based representations essential. Unlike in many publicly available 3D datasets where shapes are treated as holistic entities, real-world engineering applications therefore require structured, part-aware models to support tasks such as simulation, optimization, and interactive design refinement. For example, when optimizing the aerodynamics of a car, correctly identifying wheels is crucial for defining appropriate boundary conditions and to ensure that they stay separate from the body.
With these requirements in mind, implicit Neural Representations (INRs) (Park et al., 2019; Mescheder et al., 2019; Chen & Zhang, 2019) have become a well-established approach to providing smooth and topology-agnostic models. However, most INR-based methods employ global shape representations, making them unsuitable for these purposes. Recent INR works have explored learning part decompositions (O. et al., 2022; Chen et al., 2023), but unsupervised methods often struggle with consistency and semantic coherence—critical aspects in engineering workflows where part definitions are predefined and must remain stable across designs. Even supervised approaches (Wu et al., 2020; Deng et al., 2022; Li et al., 2024) primarily focus on generation or manipulation rather than providing a structured shape prior that enables controlled modifications or optimizations while preserving inter-part relationships.
To address these limitations, we introduce PartSDF, a supervised, part-aware implicit representation designed for composite shapes. Our approach explicitly encodes each part with a latent vector and pose parameters, ensuring independent control while maintaining global consistency. Unlike prior works that rely on complex architectures (Deng et al., 2022; O. et al., 2022; Li et al., 2024), PartSDF achieves excellent results with a simpler decoder architecture. This formulation makes it particularly effective for shape reconstruction, generation, optimization, and targeted manipulation—key requirements in engineering applications.
Our auto-decoding framework learns a structured shape prior, which can be leveraged by secondary models for various tasks. Additionally, we introduce a novel strategy for supervising parts without assuming watertight segmentation, using the global shape’s signed distance field (SDF). Each part’s SDF is an explicit output, enabling the enforcement of topological constraints during training—an aspect that, while not the primary focus of this work, may be explored in the future.
Our main contributions are as follows:
@@
  • We introduce a supervised, part-aware implicit representation that models each component independently, ensuring both expressiveness and part consistency.
@@
  • We demonstrate that our approach serves as a versatile basis for various tasks such as shape reconstruction, optimization, and manipulation, all with the same core decoder network.
@@
Through comprehensive evaluations, we show that PartSDF effectively captures the structure of composite shapes, making it well-suited for engineering applications where part-aware control is essential.
@@ 2 Related Work
Over the last several decades, the evolution of traditional CAD systems has reflected a fundamental tension between simplicity and expressiveness. Early approaches relied on basic geometric primitives such as spheres, cylinders, and NURBS surfaces (Piegl, 1991), offering mathematical precision but limited representational power. Recent years have witnessed a shift toward more sophisticated primitives, from simple cuboids (Tulsiani et al., 2017; Niu et al., 2018; Sun et al., 2019; Kluger et al., 2021) to learned deep representation, which we briefly review here. We start with generic 3D shape representations learned from data and continue with more-flexible part-based approaches.
@@ 2.1 Learned 3D Shape Representation
Learning-based methods for 3D shape representation have evolved significantly, beginning with explicit approaches such as voxel grids, point clouds, and meshes. Voxel-based methods (Wu et al., 2015, 2016; Choy et al., 2016; Dai et al., 2017) partition 3D space into grids, but are memory-intensive at finer resolutions, which can be mitigated using octrees (Riegler et al., 2017; Tatarchenko et al., 2017), but only up to a point. Point clouds reduce memory costs by representing shapes as surface points but ignore connectivity, which can compromise topological consistency (Fan et al., 2017; Yang et al., 2018; Achlioptas et al., 2018; Peng et al., 2021; Zeng et al., 2022). Meshes, though well-suited for detailed surface representation, impose a rigid topology that is difficult to modify (Groueix et al., 2018; Kanazawa et al., 2018; Wang et al., 2018; Pan & Jia, 2019).
Implicit Neural Representations (INRs) offer a flexible alternative, defining shapes as continuous functions of the 3D space that encode the surface implicitly (Park et al., 2019; Mescheder et al., 2019; Chen & Zhang, 2019; Xu et al., 2019). It can then be recovered explicitly using meshing algorithms (Lorensen & Cline, 1987; Lewiner et al., 2003; Ju et al., 2002). Even though these meshing algorithms may not be themselves differentiable, differentiability can be preserved by relying on the implicit function theorem (Guillard et al., 2024), enabling back-propagation from the explicit surface, for example when optimizing a shape to maximize its aerodynamic performance (Baqué et al., 2018). Extensions enable shape manipulation (Hao et al., 2020), point-cloud reconstruction (Peng et al., 2020), and training directly from point data (Atzmon & Lipman, 2020; Gropp et al., 2020). Newer works (Sitzmann et al., 2020; Takikawa et al., 2021) improve the accuracy further with new latent representations, such as irregular grids (Zhang et al., 2022) or unordered sets (Zhang et al., 2023).
Figure 2: PartSDF pipeline. (a) Our model’s core is a part auto-decoder $f_{\theta}$ that takes as input part latents $\mathbf{z}_{p}$ and poses expressed in terms of a quaternion $\mathbf{q}_{p}$, translation $\mathbf{t}_{p}$, and scale $\mathbf{s}_{p}$, along with the query position $\mathbf{x}\in^{3}$. It outputs signed distances $\hat{s}_{p}$ for all parts at the queried position, which may be combined into the global signed distance. (b) A secondary model may be used based on the task at hand, such as encoders to map a given modality, e.g., point clouds, to part latents and poses, or a diffusion model to generate them from noise.
@@ 2.2 Part Based Models
Early unsupervised approaches learn shape abstractions using local primitives such as cuboids (Tulsiani et al., 2017; Zou et al., 2017; Sun et al., 2019; Smirnov et al., 2020; Yang & Chen, 2021), superquadrics (Paschalidou et al., 2019, 2020), anisotropic Gaussians (Genova et al., 2019), or convexes (Deng et al., 2020), approximating complex shapes as combinations of simpler parts. More recently, complex objects are better represented by predicting and deforming primitives (Paschalidou et al., 2021; Shuai et al., 2023), creating deformable part templates (Hui et al., 2022; Chen et al., 2023), or performing part-based co-segmentation, which consistently divides shapes into parts across a dataset without relying on labeled boundaries (O. et al., 2022) or does so for only a subset of the data (Chen et al., 2019). For shape manipulation, SPAGHETTI (O. et al., 2022) utilizes a transformer to predict Gaussian parts from a global latent vector and reconstructs them into a single cohesive object. This approach supports interactive editing of user-defined parts and can be combined with SALAD (Koo et al., 2023), which employs a cascaded diffusion model for part generation. However, the parts learned by these methods are typically arbitrary and lack semantic meaning, making them unsuitable for tasks requiring specific, predefined part structures.
@@
Supervised methods leverage known part decompositions to create explicit, structured representations of composite shapes. HybridSDF (Vasu et al., 2022) mixes INRs with geometric primitives to represent and manipulate shapes, while ANISE (Petrov et al., 2023) learns to assemble implicit parts for reconstruction from images and point clouds. Instead, ProGRIP (Deng et al., 2022), PQ-Net (Wu et al., 2020), and PASTA (Li et al., 2024) focus on composite shape generation. ProGRIP relies on shape programs to produce composite INR shapes, while PQ-Net uses a GRU-based RNN (Cho et al., 2014) as an auto-encoder for sequences of parts, with latentGANs (Achlioptas et al., 2018) for generation. PASTA employs an auto-regressive transformer to predict part bounding boxes that are decoded into a single global shape, similarly to SPAGHETTI.
While effective for shape generation, these supervised methods forgo part consistency across shapes and continuous shape parametrization. Thus, there is a need for an approach that leverages part supervision to produce modular, flexible representations that can output composite shapes by parts and act as part and shape priors in backward tasks such as optimization. PartSDF aims to fulfill that need.
@@ 3 Method
We propose PartSDF, a modular, part-based representation for composite shapes, designed for flexibility across applications that require both structured part manipulation and optimization. Given a dataset of composite shapes with known part decompositions, our model learns to represent each one independently within a compact latent space, while being able to maintain consistency when deforming them.
As shown in Figure 2, our framework consists of two primary components. First, the core of our model is a composite shape decoder that outputs a signed distance function (SDF) for each part, parameterized by a latent vector and a pose (Section 3.1). It is trained in an auto-decoding fashion (Park et al., 2019) in which the latent vectors and the decoder’s weights are learned simultaneously, with supervision applied at both the global and part levels (Section 3.2). Second, our architecture supports secondary models for part encoding or generation, allowing further adaptation of part latents and poses for tasks such as shape reconstruction or generation, while maintaining the same core part decoder for efficient inference and manipulation (Section 3.3).
3.1 Composite Shape Decoder
In our composite shape representation, shown in Figure 2(a), each part is described by a latent vector $\mathbf{z}_{p}\in^{Z}$ and a pose $\mathbf{p}_{p}\in^{10}$, consisting of a rotation quaternion $\mathbf{q}_{p}\in^{4}$, translation $\mathbf{t}_{p}\in^{3}$, and scale $\mathbf{s}_{p}\in^{3}$. Given a query point $\mathbf{x}\in\mathbb{R}^{3}$, each part’s pose maps the query point into the part’s canonical space via inverse transformation $\mathcal{T}^{-1}$, resulting in the transformed query point $\hat{\mathbf{x}}_{p}=\mathcal{T}^{-1}(\mathbf{x},\mathbf{p}_{p}),\,\forall p$. The decoder $f$ then outputs the SDF for each part at this transformed point
$\hat{s}_{p}=f(\mathbf{z}_{p},\hat{\mathbf{x}}_{p}).$ (1)
@@ 3.1 Composite Shape Decoder
In PartSDF, the decoder $f$ is implemented as a multi-layer perceptron (MLP) with modifications to support part-based representation. First, we perform latent modulation (Dupont et al., 2022) by injecting the part latent vectors $\mathbf{z}_{p}$ at every fully-connected layer of the MLP. Then, we introduce a per-part learnable bias into these layers to help differentiate the individual parts. We formalize these layers $h_{\mathrm{fc}}$ as
$\mathbf{x}_{p}^{l+1}=h_{\mathrm{fc}}^{l}(\mathbf{z}_{p},\mathbf{x}_{p}^{l})=\sigma\left(\mathbf{W}^{l}\mathbf{x}_{p}^{l}+\mathbf{b}^{l}+\mathbf{W}_{z}^{l}\mathbf{z}_{p}+\mathbf{b}_{p}^{l}\right)\;,$ (2)
where $\sigma$ is an activation function, $\mathbf{x}_{p}^{l}$ denotes the features of part $p$ at layer $l$, $(\mathbf{W}^{l},\mathbf{b}^{l})$ are the layer’s parameters, $\mathbf{W}_{z}^{l}$ is the parameters of the latent modulation, and $\mathbf{b}_{p}^{l}$ the learnable bias of part $p$ at layer $l$.
This creates a part-specific MLP structure with extensive weight sharing across parts. However, it processes each part independently, which is not desirable if they are to adapt to each others when the composite shape is manipulated. To remedy this, we incorporate 1D convolutions with a kernel size of 1 between each $h_{\mathrm{fc}}$ layer. These convolutions operate across parts, treating them as the ”channel” dimension, with one filter $\mathbf{w}_{p}$ per part. Hence, for each feature dimension $d$:
${\mathbf{x}}_{p}^{l+1}[d]=\sigma\left(\sum_{i=1}^{P}\mathbf{w}_{p}^{l}[i]\cdot{\mathbf{x}}_{i}^{l}[d]+b_{p}^{l}\right),\forall p,d\;.$ (3)
Conceptually, they function as the $h_{\mathrm{fc}}$ layers, but on another input dimension: They process all parts for each feature individually (see Section B.1 for a visualization). This enables the decoder to capture inter-part dependencies and improve adaptability during part manipulation, as shown in our ablation study (Appendix E). Our final model $f_{\theta}$, with $\theta$ containing all trainable parameters, operates on all parts. We write
$\hat{\mathbf{s}}=f_{\theta}\left(\mathbf{Z},\hat{\mathbf{X}}\right)\;,$ (4)
where $\hat{\mathbf{s}}$, $\mathbf{Z}$, and $\hat{\mathbf{X}}$ are all parts’ SDF, latents, and transformed query points respectively.
3.2 Training

### v2

@@ Abstract
Accurate 3D shape representation is essential in engineering applications such as design, optimization, and simulation. In practice, engineering workflows require structured, part-based representations, as objects are inherently designed as assemblies of distinct components. However, most existing methods either model shapes holistically or decompose them without predefined part structures, limiting their applicability in real-world design tasks. We propose PartSDF, a supervised implicit representation framework that explicitly models composite shapes with independent, controllable parts while maintaining shape consistency. Thanks to its simple but innovative architecture, PartSDF outperforms both supervised and unsupervised baselines in reconstruction and generation tasks. We further demonstrate its effectiveness as a structured shape prior for engineering applications, enabling precise control over individual components while preserving overall coherence. Code available at https://github.com/cvlab-epfl/PartSDF.
@@ 1 Introduction
Engineering design is fundamentally part-driven. Whether it’s a car with distinct wheels, a chair with adaptable legs, or an industrial mixer with a removable helix, objects are designed and reasoned about in terms of their components. These parts are not just geometric regions—they carry semantics and functional roles and must satisfy a number of constraints. In real workflows, engineers rarely optimize or modify a shape as a whole. Instead, they manipulate individual parts, often under precise constraints, while expecting the overall assembly to remain coherent and valid.
Despite this, most 3D shape modeling methods relying on machine learning treat objects as indivisible units—holistic fields, point clouds, or meshes—with no modular structure. In particular, Implicit Neural Representations (INRs), first introduced in (Park et al., 2019; Mescheder et al., 2019; Chen & Zhang, 2019) and then refined in works such as Zhang et al. (2022; 2023), yield impressive fidelity and flexibility but often encode shapes as global functions with no built-in notion of parts. This makes tasks like part-specific manipulation, targeted optimization, or structured shape generation surprisingly difficult, often requiring cumbersome post-processing or bespoke architectures.
Although introducing parts into INRs might seem straightforward, it is, in fact, far from trivial. Modeling parts in isolation can easily break the continuity and alignment of the overall shape (Wu et al., 2020; Deng et al., 2022). Training such models often assumes clean, watertight geometries for each part—an assumption rarely met in real-world data. And even if each part is modeled successfully, how can we ensure that they remain responsive to one another—so that modifying one component, like enlarging the wheel of a car, naturally causes other parts, such as the car body, to adapt in response?
@@
Figure 1: Part-based implicit representation. PartSDF is a simple and modular approach to representing composite 3D shapes for many different purposes: (a) Shape generation, possibly conditioned on a part layout. (b) Part-based manipulation. (c) Part-aware optimization. In this example, the car body, shown in red, is optimized to reduce aerodynamic drag while maintaining the shape and position of the wheels fixed.
In this work, we introduce PartSDF, a simple yet effective framework for learning part-aware shape representations using signed distance functions. Each part is represented by a latent vector and a pose—capturing its geometry and spatial configuration—and decoded by a novel shared implicit model. The full shape emerges by taking the minimum across part SDFs, producing a seamless composite. This formulation is modular, supports precise part control, and naturally composes into globally coherent shapes.
To make part-based modeling practical and robust, we develop a decoder architecture that enables dynamic interaction between parts during inference. This decoder involves multiple lightweight convolutions that operate within each part and across parts respectively, enabling part-wise information exchange into a shared intermediate representation and allowing each part to adapt itself based on the geometry of others. This yields consistent deformations, without heavy computations and without resorting to complex hierarchical assemblies or attention-based systems (Deng et al., 2022; O. et al., 2022; Li et al., 2024). Complementing this, we propose a novel supervision scheme that bypasses the need for clean, watertight part meshes. Instead of requiring explicit SDF supervision for each part, we derive supervisory signals from the global shape’s SDF, applied only within regions of space closest to each part. This strategy not only aligns with the natural spatial semantics of parts but also allows effective training on noisy, incomplete, or non-watertight data—significantly broadening the method’s applicability without compromising geometric fidelity.
This results in a versatile and fully differentiable multi-part representation that is both compact and expressive. It can be applied across tasks—including shape reconstruction, part-aware generation, and constrained optimization—without modifying the core auto-decoder architecture, as illustrated in Figure˜1. Even though its architecture is simpler than that of other state-of-the-art methods, PartSDF not only improves performance on traditional benchmarks but also opens up practical capabilities that are hard to achieve with existing methods, such as optimizing a car body for drag while keeping its wheels fixed or generating diverse furniture layouts with interchangeable components.
This combination of simplicity and effectiveness is made possible by the following innovations.
@@
  • We introduce a supervised, part-aware implicit representation that models each component independently, ensuring both expressiveness and part consistency for composite objects, a key component of which is an innovative decoder that enforces consistency without undue complexity.
@@
  • We demonstrate that our approach serves as a versatile basis for various tasks such as shape reconstruction, manipulation, and optimization, all with the same core decoder network.
@@
Through comprehensive evaluations, we show that PartSDF effectively captures the structure of composite shapes, making it well-suited for engineering applications requiring part-aware representation and control.
@@ 2 Related Work
Over the last several decades, the evolution of traditional CAD systems has reflected a fundamental tension between simplicity and expressiveness. Early approaches relied on basic geometric primitives such as spheres, cylinders, and NURBS surfaces (Piegl, 1991), offering mathematical precision but limited representational power. Recent years have witnessed a shift toward more sophisticated primitives, from simple cuboids (Tulsiani et al., 2017; Niu et al., 2018; Sun et al., 2019; Kluger et al., 2021) to learned deep representation, which we briefly review here. We start with generic 3D shape representations learned from data and continue with more flexible part-based approaches.
@@ 2.1 Learned 3D Shape Representation
Learning-based methods for 3D shape representation have evolved significantly, beginning with explicit ones such as voxel grids, point clouds, and meshes. Voxel methods (Wu et al., 2015; 2016; Choy et al., 2016; Dai et al., 2017) partition 3D space into grids but are memory-intensive at finer resolutions, which can be mitigated using octrees (Riegler et al., 2017; Tatarchenko et al., 2017), but only up to a point. Point clouds reduce memory costs by representing shapes as a set of points but ignore connectivity, which can compromise topological consistency (Fan et al., 2017; Yang et al., 2018; Achlioptas et al., 2018; Peng et al., 2021; Zeng et al., 2022). Meshes, though well-suited for detailed surface representation, impose a rigid topology that is difficult to modify (Groueix et al., 2018; Kanazawa et al., 2018; Wang et al., 2018; Pan & Jia, 2019).
Implicit Neural Representations (INRs) offer a flexible alternative, defining shapes as continuous functions of the 3D space that encode the surface implicitly (Park et al., 2019; Mescheder et al., 2019; Chen & Zhang, 2019; Xu et al., 2019). It can then be recovered explicitly using meshing algorithms (Lorensen & Cline, 1987; Lewiner et al., 2003; Ju et al., 2002). Even though these meshing algorithms may not be themselves differentiable, differentiability can be preserved by relying on the implicit function theorem (Guillard et al., 2024), enabling back-propagation from the explicit surface, for example, when optimizing a shape to maximize its aerodynamic performance (Baqué et al., 2018). Extensions enable shape manipulation (Hao et al., 2020), point-cloud reconstruction (Peng et al., 2020), and training directly from point data (Atzmon & Lipman, 2020; Gropp et al., 2020). Newer works (Sitzmann et al., 2020; Takikawa et al., 2021) improve the accuracy further with new latent structures, such as grids (Yan et al., 2022; Mittal et al., 2022), irregular grids (Zhang et al., 2022) or unordered sets (Zhang et al., 2023), and using generative models within these latent spaces.
@@ 2.2 Part Based Models
Early unsupervised approaches learn shape abstractions using local primitives such as cuboids (Tulsiani et al., 2017; Zou et al., 2017; Sun et al., 2019; Smirnov et al., 2020; Yang & Chen, 2021), superquadrics (Paschalidou et al., 2019; 2020), anisotropic Gaussians (Genova et al., 2019), or convexes (Deng et al., 2020), approximating complex shapes as combinations of simpler parts. More recently, complex objects are better represented by predicting and deforming primitives (Paschalidou et al., 2021; Shuai et al., 2023), creating deformable part templates (Hui et al., 2022), or performing part-based co-segmentation, which consistently divides shapes into parts across a dataset without relying on labeled boundaries (Chen et al., 2023) or does so for only a subset of the data (Chen et al., 2019). Instead, RIM-Net (Niu et al., 2022) learns a hierarchical structure of INRs, while PartNeRF (Tertikas et al., 2023) proposes a rendering-based approach for generating part-aware editable 3D shapes. SPAGHETTI (O. et al., 2022) predicts Gaussian parts from a global latent vector and reconstructs them into a single cohesive object, supporting interactive editing of user-defined parts. It can be combined with SALAD (Koo et al., 2023), which employs a cascaded diffusion model for part generation.
However, the parts learned by these methods are typically arbitrary and lack semantic meaning, making them unsuitable for tasks requiring specific, predefined part structures.
@@
Supervised methods leverage known part decompositions to create explicit, structured representations of composite shapes. SDM-Net (Gao et al., 2019) proposes VAEs at the part and shape levels to learn deformable meshes. HybridSDF (Vasu et al., 2022) mixes INRs with geometric primitives to represent and manipulate shapes, while ANISE (Petrov et al., 2023) learns to assemble implicit parts for reconstruction from images and point clouds. Recently, DiffFacto (Nakayama et al., 2023) proposes cross-diffusion to generate and control part-based point clouds, which are, however, not suited to engineering needs such as simulations. Instead, PQ-Net (Wu et al., 2020), ProGRIP (Deng et al., 2022), and PASTA (Li et al., 2024) focus on implicit composite shape generation. PQ-Net uses a recurrent neural network (Cho et al., 2014) as an auto-encoder for sequences of parts, using a latentGAN (Achlioptas et al., 2018) for generation, while ProGRIP relies on shape programs to produce composite INR shapes. PASTA employs an autoregressive transformer to predict part bounding boxes that are decoded into a single global shape.
While effective for shape generation, these methods typically forgo part consistency across shapes, thus a continuous parametrization, and the representation of individual parts. Thus, there is a need for an approach that leverages part supervision to produce modular, flexible representations that can output composite shapes by parts and act as a prior in backward tasks such as optimization. PartSDF aims to fulfill that need.
@@ 3 Method
Figure 2: PartSDF pipeline. (a) Our model’s core is a part auto-decoder $f_{\theta}$ that takes as input part latents $\mathbf{z}_{p}$ and poses expressed in terms of a quaternion $\mathbf{q}_{p}$, translation $\mathbf{t}_{p}$, and scale $\mathbf{s}_{p}$, along with the query position $\mathbf{x}\in\mathbb{R}^{3}$. It outputs signed distances $\hat{s}_{p}$ for all parts at the queried position, which may be combined into the global signed distance. (b) A secondary model may be used based on the task at hand, such as encoders to map a given modality, e.g., point clouds, to part latents and poses, or a diffusion model to generate them from noise.
We introduce PartSDF, a modular and structured composite shape representation built on signed distance fields (SDFs). As shown in Figure˜2(a), each object is represented by a set of parts, each one being parameterized by a latent code capturing its geometry, along with pose parameters defining its spatial placement (Section˜3.1). These per-part representations are decoded into SDFs using a cross-part decoder (Section˜3.2), which supports both independent part modeling and inter-part coordination. It is trained in an auto-decoding fashion (Park et al., 2019) in which the latent vectors and the decoder’s weights are learned simultaneously, with supervision applied at both the global and part levels (Section˜3.3).
As shown in Figure˜2(b), our architecture supports secondary models for part encoding or generation, allowing further adaptation of part latents and poses for tasks such as shape reconstruction or generation, while maintaining the same core part decoder for efficient inference and manipulation (Section˜3.4).
3.1 SDF Computation
In our composite shape representation, each part is described by a latent vector $\mathbf{z}_{p}\in\mathbb{R}^{Z}$ and a pose $\mathbf{p}_{p}\in\mathbb{R}^{10}$, consisting of a rotation quaternion $\mathbf{q}_{p}\in\mathbb{R}^{4}$, translation $\mathbf{t}_{p}\in\mathbb{R}^{3}$, and scale $\mathbf{s}_{p}\in\mathbb{R}^{3}$. Given a query point $\mathbf{x}\in\mathbb{R}^{3}$, it is mapped into each part’s canonical space via inverse transformation $\mathcal{T}^{-1}$, resulting in the transformed query points
$\hat{\mathbf{x}}_{p}=\mathcal{T}^{-1}(\mathbf{x},\mathbf{p}_{p})=\mathbf{R}_{p}(\mathbf{x}-\mathbf{t}_{p})/\mathbf{s}_{p},\,\forall p\;,$ (1)
where $\mathbf{R}_{p}$ is the rotation matrix obtained from $\mathbf{q}_{p}$. Our cross-part decoder $f_{\theta}$, with $\theta$ containing all trainable parameters, operates on all parts’ latents $\mathbf{Z}$ and transformed query points $\hat{\mathbf{X}}$ to output the part SDFs as
[…]
@@ 3.1 SDF Computation
3.2 Cross-Part Auto-Decoder
To represent complex shapes in a modular yet coherent manner, we introduce an innovative decoder that explicitly handles multiple interacting parts. At each layer, the decoder maintains a multi-part feature matrix $\mathbf{X}^{l}\in\mathbb{R}^{P\times D_{l}}$, where each row corresponds to a part and each column to a shared feature dimension. This structure is maintained throughout the network, preserving an explicit notion of parts across all layers. As a result, the decoder’s output naturally yields one part per row—enabling direct, part-specific SDF predictions.
To achieve this, we design the network to alternate between two types of layers: single-part layers that update parts independently and cross-part layers that enable controlled feature sharing across parts.
Single-part layer. The first type of layer processes each part separately, allowing the network to learn the shape distribution of individual parts, denoted as $h_{\mathrm{sp}}$:
$\mathbf{x}_{p}^{l+1}=h_{\mathrm{sp}}^{l}(\mathbf{z}_{p},\mathbf{x}_{p}^{l})=\sigma\left(\mathbf{W}^{l}\mathbf{x}_{p}^{l}+\mathbf{b}^{l}+\mathbf{W}_{z}^{l}\mathbf{z}_{p}+\mathbf{b}_{p}^{l}\right)\;,$ (3)
where $\sigma$ is an activation function, $\mathbf{x}_{p}^{l}$ denotes the $p$-th row of $\mathbf{X}^{l}$, i.e., the features of part $p$ at layer $l$, $(\mathbf{W}^{l},\mathbf{b}^{l})$ are the layer’s parameters, $\mathbf{W}_{z}^{l}$ is the parameters of a latent modulation (Dupont et al., 2022), and $\mathbf{b}_{p}^{l}$ the learnable bias of part $p$ at layer $l$.
Cross-part layer. The second type of layer enables feature aggregation across parts, denoted as $h_{\mathrm{cp}}$. For each feature dimension $d$, we write
$\tilde{\mathbf{x}}_{d}^{l+1}=h_{\mathrm{cp}}^{l}(\tilde{\mathbf{x}}_{d}^{l})=\sigma\left(\tilde{\mathbf{W}}^{l}\tilde{\mathbf{x}}_{d}^{l}+\tilde{\mathbf{b}}^{l}\right)\;,$ (4)
where $\tilde{\mathbf{x}}^{l}$ denotes the $d$-th column of $\mathbf{X}^{l}$, i.e., the $d$-th feature dimension across all parts, and $(\tilde{\mathbf{W}}^{l},\tilde{\mathbf{b}}^{l})$ are the layer’s parameters. This layer enables the decoder to capture inter-part dependencies and improve adaptability during part manipulation with as little as $P^{2}+P$ added trainable parameters per such layer, with $P$ the number of parts, as shown in our ablation study of Section˜4.5. We also add a skip connection to ensure these layers don’t mix part identities.
By stacking alternating $h_{\mathrm{sp}}$ and $h_{\mathrm{cp}}$ layers (see Figure˜4), which can be efficiently implemented using convolutions, the decoder enables each part to specialize while remaining aware of its context—producing modular, adaptive, and coherent shape representations. We provide network details in Section˜B.1.
Figure 3: Cross-part adaptation in PartSDF. Our decoder alternates between updating each part independently and sharing information across parts, allowing them to adapt to one another while preserving modularity. This is implemented through a sequence of lightweight convolutions applied along rows and columns of the part feature matrix.